# HIF: The hypergraph interchange format for higher-order networks

### v1

@@ III.2 Checking HIF compliance
A central contribution of the HIF project is a JSON schema allowing simple validation of datasets against the standard. There are simple JSON validators in several different languages and we illustrate simple ways to validate datasets against the HIF Standard in Fig. 2.
@@
We emphasize that the validating the schema should be the first step in the implementation of an independent software packages which utilize this standard are ultimately responsible for reading and writing datasets to and from the appropriate network representation. For example, the standard uses the ”asc” keyword to specify that every interaction present in the hypergraph is downward closed, but checking the dataset against the schema doesn’t explicitly check this. In addition, best practices are that only the maximal simplex faces and subfaces with corresponding attributes are stored for simplicial complexes, but this informal standard is not enforced by the schema. Lastly, while end users are able to, in principle, specify the direction keyword for network datasets even if the network-type keyword is not directed and to omit the direction keyword even when network-type is directed, for simplicity, we simply say that direction field is optional.
III.3 Updating the schema
@@
    Decide on the new version number based on the changes made since the last version and the Semantic Versioning guidelines [48].
@@
The hif_ schema.json schema will always have version ”latest” and all unit tests are based on this schema.
IV Integration with software libraries
@@
A strength of the HIF standard is its integration with five prominent software libraries for higher-order network analysis: the Hypergraph Analysis Toolbox (HAT) [24], Hypergraphx [26], HyperNetX [27], SimpleHypergraphs.jl [29], and XGI [31]. Here we detail the ways in which each library supports the HIF standard as well as brief descriptions of their use.
IV.1 Hypergraphx
@@
Hypergraphx (HGX) is an open-source Python library designed for the analysis of complex systems characterized by higher-order interactions. HGX offers an extensive collection of tools and algorithms for constructing, manipulating, analyzing, and visualizing hypergraphs, enabling users to capture the rich structure and study the dynamics of real-world systems beyond pairwise relationships. HGX enables seamless storage and conversion of higher-order data across multiple representations, including hypergraphs, simplicial complexes, bipartite networks, line graphs and clique expansions, all while supporting hyperedges with weights, direction, sign, temporal dynamics and multiplex layers. The library provides tools for basic statistical characterizations (e.g., hyperedge size distribution), higher-order centrality, motif analysis (with scalable sampling), community detection, significant interaction filtering, and synthetic hypergraph generation. HGX also supports the simulation of dynamical processes and offers advanced visualization functionalities for exploring the structure of real-world systems.
@@
On the data and I/O side, HGX is accompanied by HGX-data, a collection of ready-to-use datasets for higher-order network analysis in Python, including collaboration, face-to-face, biological data, and more. HGX offers dedicated input and output functions, including a binary format to enable fast and efficient I/O over large-scale higher-order datasets. Hypergraphx has integrated HIF support through the hypergraphx.readwrite.read_ hif function, allowing users to load hypergraphs directly from HIF-compliant JSON files and automatically handling the conversion to HGX’s internal data structures. Similarly, the hypergraphx.readwrite.write_ hif function enables users to export an HGX Hypergraph object to the HIF format, preserving both the structural information and any associated node, edge, or incidence attributes. This interoperability allows users to seamlessly leverage HGX functionalities while moving back and forth across different libraries. In the future, HIF will accommodate additional hypergraph representations available in Hypergraphx, including multiplex hypergraphs. Moreover, HGX-data is planned to be fully HIF-compatible in order to be accessed by the whole higher-order software ecosystem.
@@
An example demonstrating how to load a hypergraph from a HIF file, leverage HGX functionalities such as motif analysis, and export the data is provided in Fig. 3.
@@
IV.2 HyperNetX
HyperNetX (HNX) is a python package focused on developing tools for the analysis and visualization of hypergraphs. HNX provides a user-friendly interface for creating, manipulating, and studying hypergraphs, with an emphasis on visualization capabilities that capture higher-order relationships. Some of the key features of HNX include:
  • A suite of analytical functions, including degree sequences, connectivity measures, and centrality metrics.
@@ IV.2 HyperNetX
HyperNetX has integrated HIF through the hnx.read_ hif and hnx.to_ hif functions: hnx.read_ hif loads hypergraphs directly from HIF-compliant JSON files, handling the conversion to HNX’s internal data structure and hnx.to_ hif exports an HNX hypergraph object to the HIF format, preserving both the structural information and any associated node, edge, or incidence attributes. An example demonstrating how to load and save an HNX hypergraph is shown in Fig. 4 where the filename is provided. Alternatively, the JSON object can be loaded first and then the JSON object converted to an HNX hypergraph can be done.
@@
IV.3 SimpleHypergraphs.jl
SimpleHypergraphs.jl is a Julia package designed to efficiently work with hypergraphs [29]. Its key capabilities include (1) creating hypergraphs from incidence matrices, random hypergraph models, a Graphs.jl graph, or external data; (2) manipulating hypergraphs, (3) analyzing hypergraphs with fast algorithms for computing, for example, connected components or modularity based clustering; (4) visualizing hypergraphs; (5) integration with the Graphs.jl Julia package via bipartite graphs or a hypergraph projection; and (6) and file I/O. SimpleHypergraphs.jl is designed to support (1) two-way of storage of node and hyperedge incidence information, which enables efficient traversal of the hypergraph in both directions — from nodes to hyperedges and vice versa; (2) hyperedge-dependent node weights, i.e., incidence weights; (3) node and hyperedge metadata; (4) compatibility with Julia’s AbstractMatrix interface, enabling efficient manipulation of hypergraphs using standard Julia matrix operations; and (5) lazy representation of hypergraph views which are compatible with the AbstractGraph interface of the Graphs.jl library, enabling seamless integration and direct usage of Graphs.jl algorithms. Figure 5 presents an example of SimpleHypergraphs.jl library usage with HIF.
@@
IV.4 Hypergraph Analysis Toolbox
The Hypergraph Analysis Toolbox (HAT) [24] is a general-purpose software for constructing, visualizing, and analyzing hypergraphs and higher-order structures, with a focus on their structure and dynamics. HAT implements tensor based methods for the analysis of higher-order observability/controllability [49, 50], coupling [51], and structural properties suitable for analysis of a wide range of data [52, 53]. It provides HIF-compliant read and write capabilities in the HAT.Hypergraphs module. Figure 6 presents a sample using the HAT library. The Hypergraph.to_ hif method converts the hypergraph object back to a Python dictionary according to the HIF schema, which can be saved to a JSON file.
IV.5 XGI
The CompleX Group Interactions (XGI) [31] software package provides support for hypergraphs, directed hypergraphs, and simplicial complexes. It provides a comprehensive suite of analysis and visualization tools, including algorithms, generative higher-order network models, linear algebra representations, conversions between many different data structures, and an integrated statistical interface. Integrated with this software environment is the XGI-DATA repository, which hosts many higher-order datasets. XGI provides not only visualization and analysis tools for higher-order networks but the ability to read and write numerous different file formats such as bipartite edge lists, hyperedge lists, and incidence matrices.
XGI currently supports the HIF standard in two ways: first, it provides HIF-compliant read and write capabilities and second, it provides HIF-compliant datasets in the XGI-DATA collection. XGI has implemented an hif sub-module in its readwrite module, providing two methods: read_ hif and write_ hif. These methods allow XGI to store undirected hypergraphs, directed hypergraphs, and simplicial complexes according to the HIF standard. When storing directed hypergraphs, XGI makes use of the direction field in the incidence records. When storing simplicial complexes, XGI only stores the maximal faces and hyperedges with associated attributes, with the expectation that the asc keyword will indicate that XGI should enforce downward closure when reading in the file. The HIF standard is robust to the future development of XGI; for example, it implements features not yet supported by XGI. One example is edge-dependent node attributes, also known as incidence attributes, which capture properties of physical systems such as stoichiometric coefficients in chemical reactions and interaction-dependent social roles.
@@ V.1 Dataset
We use a publication dataset for our case study because of its rich attributes and tractable size. This dataset consists of open-source publications with the keyword ”Hypergraph” and was collected from ArXiv, BioRxiv, the DBLP computer science bibliography, and the U.S. Office of Scientific and Technical Information (OSTI). The resulting hypergraph is composed of hyperedges, which represent scientific publications, and nodes, which represent authors. Each hyperedge (publication) has a number of attributes including funding agencies, abstract, publication date, keyword tags, and the source, while each node (author) has an associated institution have the attributes of institutions (There are no incidence attributes). The hypergraph has 1,960 nodes (authors) and 533 edges (publications). The following analysis of this data using each of the hypergraph packages is provided in the notebook HIF-demo.ipynb in the tutorials folder of the HIF GitHub page  [^2] See https://github.com/pszufe/HIF-standard.. We briefly describe examples of the types of analysis that one can perform on this dataset for each software package.
@@
Figure 9: The Hypergraph Interchange Format enables interoperability between many popular higher-order network analysis libraries, as seen in this case study which analyzes a publication dataset. The notebook that creates each visualization can be found in the HIF-demo.ipynb file. In the center, we display the HIF logo. In panel (a), we use HAT to compute the nonlinear eigenvector centralities of the nodes and hyperedges. In panel (b), we use Hypergraphx to visualize the largest connected component of the hypergraph (i), with node colors indicating the community structure. We also visualize the most frequent patterns of group interactions involving three (ii) and four nodes (iii). In panel (c), we use HyperNetX to solely examine the largest component of the hypergraph, compute the closeness centrality for each hyperedge, and color each hyperedge according to its centrality as we can see from the color bar. In panel (d), we use SimpleHypergraphs.jl to find the largest component, compute modularity-based community detection, and to visualize the pairwise network projected from the hypergraph, where each hyperedge becomes a clique. In panel (e), we use XGI to compute nodal statistics using the statistics interface; in the visualization, the node size is scaled by the nodal degree and the node color corresponds to the clique eigenvector centrality (CEC) [55]. Similarly, in the top inset figure, we see the degree (solid line) and average neighbor degree (dashed line), while in the bottom inset figure, we see the CEC (solid line) and the H-eigenvector centrality (HEC) [55, 56] (dashed line). Note that the node labels are sorted in descending order according to the degree (top) and CEC (bottom).
V.1.1 Hypergraph Analysis Toolbox
For HAT we demonstrate the usefulness of analyzing the publications hypergraph using the eigenvector centrality measure. This allows us to analyze node and edge centrality. In 9(a), we compute the centralities of each node and edge.
V.1.2 Hypergraphx
@@
To demonstrate the capabilities of Hypergraphx, we consider our example publications dataset and perform an exploratory analysis combining community detection, motifs, and visualization. Specifically, in Fig. 9(b), we apply (overlapping) community detection to the largest connected component, identifying groups of closely related nodes [57], and visualize the structure of the resulting hypergraph. We further use Hypergraphx to mine the frequency of subhypergraph patterns of sizes 3 and 4 [58], describing the hypergraph of publications at its microscale and providing insights into the structural building blocks of higher-order interactions present in the data [59]. This use case illustrates how Hypergraphx can serve as a key component of an integrated workflow within the higher-order network ecosystem, enabling complex tasks such as motif analysis and community detection to be performed efficiently and in combination with other tools.
V.1.3 HyperNetX

### v2

@@ 3.2 Checking HIF compliance
A central contribution of the HIF project is a JSON schema allowing simple validation of datasets against the standard. There are simple JSON validators in several different languages and we illustrate simple ways to validate datasets against the HIF Standard in Figure 2.
Python
import fastjsonschema
import json
import requests
schema = requests.get(url).json()
validator = fastjsonschema.compile(schema)
hiftext = json.load(open(filename,’r’))
try:
validator(hiftext)
print(‘‘HIF-Compliant JSON.’’)
except Exception as e:
print(f‘‘Invalid JSON: {e}’’)
R
library(jsonvalidate)
library(jsonlite)
schema <- paste(readLines(url, warn = FALSE))
validator <- json_validator(schema)
if (validator(filepath)) {
print(‘‘HIF-Compliant JSON.’’)
} else {
print(‘‘Invalid JSON.’’)
Julia
using HTTP
using JSON3
using JSONSchema
schema = String(HTTP.get(url).body)
validator = Schema(schema)
hiftext = JSON3.read(filepath)
result = JSONSchema.validate(validator, hiftext)
println(
result === nothing ?
‘‘HIF-Compliant JSON.’’ :
‘‘Invalid JSON: ’’ * ‘‘\$result’’
@@
We emphasize that the validating the schema should be the first step in the implementation of an independent software packages which utilize this standard are ultimately responsible for reading and writing datasets to and from the appropriate network representation. For example, the standard uses the “asc” keyword to specify that every interaction present in the hypergraph is downward closed, but checking the dataset against the schema doesn’t explicitly check this. In addition, best practices are that only the maximal simplex faces and subfaces with corresponding attributes are stored for abstract simplicial complexes, but this informal standard is not enforced by the schema. Lastly, while end users are able to, in principle, specify the direction keyword for network datasets even if the network-type keyword is not directed and to omit the direction keyword even when network-type is directed, for simplicity, we simply say that direction field is optional.
3.3 Updating the schema
@@
    Decide on the new version number based on the changes made since the last version and the Semantic Versioning guidelines preston-werner_semantic_2023.
@@
The hif_ schema.json schema will always have version “latest” and all unit tests are based on this schema.
4 Integration with software libraries
@@
A strength of the HIF standard is its integration with five prominent software libraries for higher-order network analysis: the Hypergraph Analysis Toolbox (HAT) pickard_hat_2023, Hypergraphx lotito_hypergraphx_2023, HyperNetX praggastis_hypernetx_2024, SimpleHypergraphs.jl spagnuolo_analyzing_2020, and XGI landry_xgi_2023. Here we detail the ways in which each library supports the HIF standard as well as brief descriptions of their use.
4.1 Hypergraph Analysis Toolbox
The Hypergraph Analysis Toolbox (HAT) [^3] https://hypergraph-analysis-toolbox.readthedocs.io pickard_hat_2023 is a general-purpose software for constructing, visualizing, and analyzing hypergraphs and higher-order structures, with a focus on their structure and dynamics. HAT implements tensor based methods for the analysis of higher-order observability/controllability pickard_observability_2023; pickard_geometric_2024, coupling pickard_kronecker_2024, and structural properties suitable for analysis of a wide range of data chen_tensor_2020; dotson_deciphering_2022. It provides HIF-compliant read and write capabilities in the HAT.Hypergraphs module. Figure 3 presents a sample using the HAT library. The Hypergraph.to_ hif method converts the hypergraph object back to a Python dictionary according to the HIF schema, which can be saved to a JSON file.
import json
from HAT import Hypergraph
with open(‘‘example_hif.json’’) as file:
hif_data = json.load(file)
HG = Hypergraph.from_hif(hif_data)
hif_output = HG.to_hif()
4.2 Hypergraphx
@@
Hypergraphx (HGX) [^4] https://github.com/HGX-Team/hypergraphx is an open-source Python library designed for the analysis of complex systems characterized by higher-order interactions. HGX offers an extensive collection of tools and algorithms for constructing, manipulating, analyzing, and visualizing hypergraphs, enabling users to capture the rich structure and study the dynamics of real-world systems beyond pairwise relationships. HGX enables seamless storage and conversion of higher-order data across multiple representations, including hypergraphs, abstract simplicial complexes, bipartite graphs, line graphs and clique expansions, all while supporting hyperedges with weights, direction, sign, temporal dynamics and multiplexity. The library provides tools for basic statistical characterizations (e.g., hyperedge size distribution), higher-order centrality, motif analysis (with scalable sampling), community detection, filtering, and synthetic hypergraph generation. HGX also supports the simulation of dynamical processes and offers advanced visualization functionalities for exploring the structure of real-world systems.
@@
On the data and I/O side, HGX is accompanied by HGX-data, a collection of ready-to-use datasets for higher-order network analysis in Python, including collaboration, face-to-face, biological data, and more. HGX offers dedicated input and output functions, including a binary format to enable fast and efficient I/O over large-scale higher-order datasets. Hypergraphx has integrated HIF support through the read_ hif function, allowing users to load hypergraphs directly from HIF-compliant JSON files and automatically handling the conversion to HGX’s internal data structures. Similarly, the write_ hif function enables users to export an HGX Hypergraph object to the HIF format, preserving both the structural information and any associated node, edge, or incidence attributes. This interoperability allows users to seamlessly leverage HGX functionalities while moving back and forth across different libraries. In the future, HIF will accommodate additional hypergraph representations available in Hypergraphx, including multiplex hypergraphs lotito_multiplex_2024. Moreover, HGX-data is planned to be fully HIF-compatible in order to be accessed by the whole higher-order software ecosystem.
@@
An example demonstrating how to load a hypergraph from a HIF file, leverage HGX functionalities such as motif analysis, and export the data is provided in Figure 4.
import hypergraphx as hgx
# Load HIF hypergraph data to HGX object
from hypergraphx.readwrite import read_hif
HG = read_hif(filename=‘‘example_hif.json’’)
# Compute and evaluate motifs
from hypergraphx.motifs import compute_motifs
m = compute_motifs(HG, runs_config_model=20)
# Save HGX hypergraph object to HIF file
from hypergraphx.readwrite import write_hif
write_hif(HG, filename=‘‘example_hif.json’’)
@@
4.3 HyperNetX
HyperNetX (HNX) (praggastis_hypernetx_2024) [^5] https://github.com/pnnl/HyperNetX is a python package focused on developing tools for the analysis and visualization of hypergraphs. HNX provides a user-friendly interface for creating, manipulating, and studying hypergraphs, with an emphasis on visualization capabilities that capture higher-order relationships. Some of the key features of HNX include:
  • A suite of network analytical functions, including line graphs, degree sequences, connectivity measures, and centrality metrics specifically modified to be appropriate for hypernetworks (aksoy_hypernetwork_2020).
  • Support for topological measures of hypergraphs, including simplicial and zigzag homology of temporal hypergraphs (myers_topological_2023).
@@ 4.3 HyperNetX
  • Advanced analytical capabilities through supplementary modules including generative models, Laplacian clustering, and hypergraph modularity.
HyperNetX has integrated HIF through the read_ hif and to_ hif functions: read_ hif loads hypergraphs directly from HIF-compliant JSON files, handling the conversion to HNX’s internal data structure and to_ hif exports an HNX hypergraph object to the HIF format, preserving both the structural information and any associated node, edge, or incidence attributes. An example demonstrating how to load and save an HNX hypergraph is shown in Figure 5 where the filename is provided. Alternatively, the JSON object can be loaded first and then the JSON object converted to an HNX hypergraph can be done.
import json
import hypernetx as hnx
# Load HIF hypergraph data to hnx object
HG = hnx.from_hif(filename = ‘‘example_hif.json’’)
# Save HIF data to file
hnx.to_hif(HG, filename=‘‘example_hif.json’’)
@@
4.4 SimpleHypergraphs.jl
SimpleHypergraphs.jl [^6] https://github.com/pszufe/SimpleHypergraphs.jl is a Julia package designed to efficiently work with hypergraphs spagnuolo_analyzing_2020. Its key capabilities include (1) creating hypergraphs from incidence matrices, random hypergraph models, a Graphs.jl graph, or external data; (2) manipulating hypergraphs, (3) analyzing hypergraphs with fast algorithms for computing, for example, connected components or modularity based clustering; (4) visualizing hypergraphs; (5) integration with the Graphs.jl Julia package via bipartite graphs or a hypergraph projection; and (6) and file I/O. SimpleHypergraphs.jl is designed to support (1) two-way of storage of node and hyperedge incidence information, which enables efficient traversal of the hypergraph in both directions — from nodes to hyperedges and vice versa; (2) hyperedge-dependent node weights, i.e., incidence weights; (3) node and hyperedge metadata; (4) compatibility with Julia’s AbstractMatrix interface, enabling efficient manipulation of hypergraphs using standard Julia matrix operations; and (5) lazy representation of hypergraph views which are compatible with the AbstractGraph interface of the Graphs.jl library, enabling seamless integration and direct usage of Graphs.jl algorithms. Figure 6 presents an example of SimpleHypergraphs.jl library usage with HIF.
using SimpleHypergraphs
hg = Hypergraph{Int, String, String}(
[1 nothing 2 nothing;
3 1 nothing 4])
set_vertex_meta!(hg, ‘‘vertex 1’’, 1)
set_hyperedge_meta!(hg, ‘‘h-edge 2’’, 2)
g_save(‘‘hg_hif.json’’, hg; format=HIF_Format())
loaded_hg = hg_load(
‘‘hg_hif.json’’;
format=HIF_Format(),
HType=Hypergraph,
T=Int, V=String, E=String
@@
4.5 XGI
The CompleX Group Interactions (XGI) [^7] https://xgi.readthedocs.io landry_xgi_2023 software package provides support for hypergraphs, directed hypergraphs, and abstract simplicial complexes. It provides a comprehensive suite of analysis and visualization tools, including algorithms, generative higher-order network models, linear algebra representations, conversions between many different data structures, and an integrated statistical interface. Integrated with this software environment is the XGI-DATA repository, which hosts many higher-order datasets. XGI provides not only visualization and analysis tools for higher-order networks but the ability to read and write numerous different file formats such as bipartite edge lists, hyperedge lists, and incidence matrices.
XGI currently supports the HIF standard in two ways: first, it provides HIF-compliant read and write capabilities and second, it provides HIF-compliant datasets in the XGI-DATA collection. XGI has implemented an hif sub-module in its readwrite module, providing two methods: read_ hif and write_ hif. These methods allow XGI to store undirected hypergraphs, directed hypergraphs, and abstract simplicial complexes according to the HIF standard. When storing directed hypergraphs, XGI makes use of the direction field in the incidence records. When storing abstract simplicial complexes, XGI only stores the maximal faces and hyperedges with associated attributes, with the expectation that the asc keyword will indicate that XGI should enforce downward closure when reading in the file. An example demonstrating how to create a hypergraph, save the hypergraph as an HIF file, and then loading that file is shown in Figure 7.
import xgi
H = xgi.Hypergraph(
[[1, 2, 3], [2, 3, 4], [1, 4]]
example_file = ‘‘example_hif.json’’
xgi.write_hif(H, example_file)
H = xgi.read_hif(example_file, nodetype=str)
H.edges.members()
[{‘‘1’’, ‘‘2’’, ‘‘3’’}, {‘‘2’’, ‘‘3’’, ‘‘4’’}, {‘‘1’’, ‘‘4’’}]
@@ 5.1 Dataset
We use a publication dataset for our case study because of its rich attributes and tractable size. This dataset consists of open-source publications with the keyword “Hypergraph” and was collected from ArXiv, BioRxiv, the DBLP computer science bibliography, and the U.S. Office of Scientific and Technical Information (OSTI). The resulting hypergraph is composed of hyperedges, which represent scientific publications, and nodes, which represent authors and can be seen in Figure 8. Each hyperedge (publication) has a number of attributes including funding agencies, abstract, publication date, keyword tags, and the source, while each node (author) has an associated institution have the attributes of institutions (There are no incidence attributes). The hypergraph has 1,960 nodes (authors) and 533 edges (publications). A detailed description and the following analysis of this dataset using each of the hypergraph packages is provided in the notebook HIF-demo.ipynb in the tutorials folder of the HIF GitHub page  [^8] See https://github.com/pszufe/HIF-standard.. We briefly describe examples of the types of analysis that one can perform on this dataset for each software package.
@@
Figure 9: The Hypergraph Interchange Format enables interoperability between many popular higher-order network analysis libraries, as seen in this case study which analyzes a publication dataset. All code to reproduce this figure is available on the HIF-standard repository (See https://github.com/pszufe/HIF-standard). The notebook that creates each visualization can be found in the HIF-demo.ipynb file. In the center, we display the HIF logo. In panel (a), we use HAT to compute the nonlinear eigenvector centralities of the nodes and hyperedges. In panel (b), we use Hypergraphx to visualize the largest connected component of the hypergraph (i), with node colors indicating the community structure. We also visualize the most frequent patterns of group interactions involving three (ii) and four nodes (iii). In panel (c), we use HyperNetX to solely examine the largest component of the hypergraph, compute the closeness centrality for each hyperedge, and color each hyperedge according to its centrality as we can see from the color bar. In panel (d), we use SimpleHypergraphs.jl to find the largest component, compute modularity-based community detection, and to visualize the network projected from the hypergraph, where each hyperedge becomes a clique. In panel (e), we use XGI to compute nodal statistics using the statistics interface; in the visualization, the node size is scaled by the nodal degree and the node color corresponds to the clique eigenvector centrality (CEC) benson_three_2019. Similarly, in the top inset figure, we see the degree (solid line) and average neighbor degree (dashed line), while in the bottom inset figure, we see the CEC (solid line) and the H-eigenvector centrality (HEC) benson_three_2019; aksoy_scalable_2024 (dashed line). Note that the node labels are sorted in descending order according to the degree (top) and CEC (bottom).
5.1.1 Hypergraph Analysis Toolbox
For HAT we demonstrate analyzing the publications hypergraph using the eigenvector centrality measure. This allows us to analyze node and edge centrality indicating the contributions of each author and publication. In Figure 9(a), the incidence matrix of the largest connected component of the hypergraph is shown, along with the node and hyperedge centrality measures. Each row of the incidence matrix indicates an author (node), and the columns of the matrix correspond to publications (hyperedges). HAT computes the centralities of the nodes and hyperedges, displaying them along the corresponding modes of the incidence matrix.
5.1.2 Hypergraphx
@@
To demonstrate the capabilities of Hypergraphx, we consider our example publications dataset and perform an exploratory analysis combining community detection, motifs, and visualization. Specifically, in Figure 9(b), we apply (overlapping) community detection to the largest connected component, identifying groups of closely related nodes contisciani_inference_2022, and visualize the structure of the resulting hypergraph. We further use Hypergraphx to mine the frequency of subhypergraph patterns of sizes 3 and 4 lotito_exact_2024, describing the hypergraph of publications at its microscale and providing insights into the structural building blocks of higher-order interactions present in the data lotito_higherorder_2022a. This use case illustrates how Hypergraphx can serve as a key component of an integrated workflow within the higher-order network ecosystem, enabling complex tasks such as motif analysis and community detection to be performed efficiently and in combination with other tools.
5.1.3 HyperNetX